# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v1

@@ Appendix B Understanding the impact of many observations
Motivated by the need for practical projections for third-generation instruments and the many-detection era, in this appendix we provide a simple ansatz allowing simple estimates showing how statistical measurement errors on the nuclear equation of state will improve as more observations accumulate. Our argument has five key ingredients. First, we will assume all binaries are equal mass. Second, following Zhao and Lattimer 2018 [145] we will assume $\Lambda(m)=\Lambda_{*}(m/m_{*})^{-p}$ with $p=6$, where $\Lambda_{*}$ (the equation of state parameter to be measured with many observations) is the value of the tidal deformability at a reference mass $m_{*}$ (unless otherwise noted, assumed to be at or close to $m_{*}=1.4M_{\odot}$). Third, following the discussion in Section III.A or Ref. [106], we will argue the marginal one-dimensional measurement accuracy for the tidal deformability scales as $\sigma_{\Lambda}=\sigma_{*}(m/m_{*})^{-1}$ versus binary mass, adopting a fiducial uncertainty for the faintest source used in our analysis of $\sigma_{*}\simeq 200$; see also, e.g., Fig. 1 in [169]. Fourth, we will assume that each independent observation’s marginal tidal deformability has a log-likelihood of the form
$\Gamma_{**,1}=\Big{(}\frac{\rho}{8}\Big{)}^{2}\sigma_{\lambda}^{-2}\Big{(}\frac{d\Lambda}{d\Lambda_{*}}\Big{)}^{2}=\Big{(}\frac{\rho}{8}\Big{)}^{2}\sigma_{*}^{-2}\Big{(}\frac{m}{m_{*}}\Big{)}^{2-2p}$ (14)
Fifth and last, we will assume the signal amplitude $\rho$ scales with mass and luminosity distance as $(m/m_{*})^{5/6}d_{L,*}/d_{L}$ times geometrical angular factors of order unity, i.e.
@@
where in the second line we have replaced the discrete sum by an ensemble average ($\sum_{k}\cdot=N{\left\langle\cdot\right\rangle}$) and for convenience in our ansatz we assume the angular, distance, and mass distributions of the $N$ detected sources are independent. The average ${\left\langle w^{2}\right\rangle}$ of the geometrical factors depends on the assumed network beampattern (and, for realistic heterogeneous networks, their duty cycle), but is well known to be ${\left\langle w^{2}\right\rangle}=2/5$ for the single-interferometer scenario; see, e.g., Finn and Chernoff [172], Cutler and Flanagan [173], or O’Shaughnessy et al. [174] for discussion. Unless otherwise noted, for the purposes of discussion we will assume the the single-interferometer beampattern factor; a uniform NS mass distribution between $1M_{\odot}$ and $2M_{\odot}$; and use ${\left\langle\Big{(}\frac{d_{L,*}}{d_{L}}\Big{)}^{2}\right\rangle}\simeq 1$ as appropriate to a third-generation network. A larger value of this expression corresponds to smaller marginal measurement accuracy $\sigma_{\Lambda_{*,N}}\equiv 1/\sqrt{\Gamma_{**,N}}$ on the fiducial tidal deformability $\Lambda_{*}$.
@@
The expression above [Eq. (16)] lets us quickly assess the relative importance of different contributions to the integral. As a concrete example, the relative contribution of sources at different distances scales as the relative contributions to ${\left\langle 1/d_{L}^{2}\right\rangle}$, which under the conventional assumption of an event rate comparable to comoving volume has a critical contribution from the weakest sources available to the population average. To be concrete, for a Euclidean distribution of sources out to the maximum detectable distance $d_{L,*}$, within our ansatz we find ${\left\langle(d_{L,*}/d_{L})^{2}\right\rangle}=3$. Weak sources play an even more important role if sources occur more frequently at higher redshift, as expected given the increasing star formation rate with redshift. In the extreme limit of strong redshift dependence and cosmological factors, we expect ${\left\langle(d_{L,*}/d_{L})^{2}\right\rangle}\simeq 1$. The primary limitation of this interpretation is the implicit assumption of a Gaussian joint likelihood, which only applies for loud signals (e.g., comparable to GW170817). Even after restricting our calculation to sources louder than $\rho_{c}$, the detection rate of pertinent sources is only reduced by of order $(8/\rho_{c})^{3}$. For third-generation instruments with $O(10^{5})$ sources, many thousands must still make important contributions, even for the most conservative choices for $\rho_{c}$. Moreover, once in the Gaussian-likelihood regime, our calculation scales much more slowly than the naive Euclidean factor, only as $N{\left\langle(d_{L,*}/d_{L})^{2}\right\rangle}$.
@@
Our expression provides an even stronger conclusion about the relative importance of different parts of the mass distribution: the lowest mass binaries overwhelmingly dominate the mass average ${\left\langle m^{11/3-2p}\right\rangle}\simeq{\left\langle m^{-8}\right\rangle}$ (more precisely = $\frac{1-2^{-(8-2/3)}}{9-2/3}$ for the average for a uniform mass distribution from 1 M to 2 M). In retrospect, the lowest mass binaries must dominate: $\Lambda(m)$ is vastly larger at the lowest allowed masses. If any binaries exist at these masses, their observations will dominate our constraints on the equation of state, because their large tidal effects will be easily differentiated from the noise.
@@
Our simple estimate is comparable to (but calibrated above to be slightly more optimistic than) detailed calculations produced in other studies [175, 103, 169]; for example, [103] argue the 90% credible interval is $\Delta\Lambda_{*}\simeq 910/\sqrt{N}$ and thus that the comparable $68\%$ credible interval would be $520/\sqrt{N}$. The aforementioned measurement accuracy on $\Lambda_{*}$ implies a relative measurement accuracy on $\ln R_{*}$ of $\Lambda_{*}^{-1}|d\ln R_{*}/d\ln\Lambda_{*}|/\sqrt{\Gamma_{**,N}}$. To evaluate this expression and put the result in context, we’ll assume for the purposes of illustration the MPA1 equation of state, which produces for $m_{*}=1.4M_{\odot}$ the fiducial pair $(\Lambda_{*},R_{*}/{\rm km})=(513,12.5)$. For simplicity again following Zhao and Lattimer [145], we can evaluate this change-of-variable expression by adopting the same scaling $\Lambda=a(m/R)^{-p}$ assumed above for some constant $a\simeq 0.0085$, so $d\ln R_{*}/d\ln\Lambda_{*}=p$. We therefore estimate the measurement accuracy $\sigma_{R_{*},N}$ in $R_{*}$ produced by stacking $N$ measurements to be
@@
Our analysis above suggests important trends in tension with previous preliminary investigations. Many groups over several years have argued that only the loudest handful of observations will dominate overall EoS inference [167, 176]. Our analytic argument suggests a comparable and potentially even dominant contribution from more numerous fainter events.
In the discussion above, we have only provided estimates for the statistical accuracy $\sigma_{\Lambda_{*},N}$ and $\sigma_{R_{*},N}$ for the tidal deformability and radius of a fiducial neutron star, respectively, after combining $N$ events. Given proposed network sensitivity models and reference, these expressions can be immediately translated to the operating time needed to constrain these quantities to any target precision. For example, in terms of the proposed detection rate of useful events (scaled to $10^{5}/{\rm yr}$, assuming all events contribute) we find for the simple estimates provided above that a target accuracy of $\Delta R_{*}=10{\rm m}$ requires an observing time scaling with event rate and fiducial neutron star radius as
@@ Appendix C More alternate perspectives on GW170817
Figure 17 shows the results of an analysis that incorporates an alternative interpretation of GW170817 performed by Lange et al. (in prep), obtained by comparing the TEOBResumS waveform model to the data. The differences in interpretation reported there propagate to modest differences in the mass-radius posterior distribution.
@@ Appendix D Hyperpipeline examples
In the text, we have presented the end results of our iterative strategy for hyperparameter inference. This approach as used for GW source parameter inference with RIFT, this appendix provides a few illustrations showing how our iterative strategy works within the context of EoS hyperparameter inference.
As a concrete example, Figure 18 shows 90% credible intervals estimated at each stage of one of our analyses, along with (in color) the marginal likelihoods accumulated during the analysis.
Initial grid is an initial guess made during this work, however the pipeline starts exploring regions to obtain a more complete picture of the posterior. By the 5th, 6th, and 7th iteration the posterior estimation does not change significantly, leading us to a converged posterior distribution.
@@ Appendix E NICER PSR-J0437
One of the NICER collaboration groups announced the observational constraints for PSR-J0437 [177]. The pulsar has a strong a-priori mass constraint from radio observations [178]. The full posterior data release has not occurred, however, here we apply this pulsar’s separate 90% credible intervals for its mass and radius and can construct an approximate Gaussian posterior to explore its effect on the EoS. The reported Radius is $R=11.36^{+0.95}_{-0.63}~{}km$ and $M=1.418\pm 0.037M_{\odot}$. For this preliminary application, we employ the constraints as a Gaussian with $R=11.36\pm 0.95~{}km$ and $M=1.418\pm 0.037M_{\odot}$.
Figure 19 displays the posterior Mass-Radius 90% credible interval shown earlier in Fig. 11, and a new posterior obtained upon applying the mass-radius constraint for PSR-J0437.

### v2

@@ Appendix B Understanding the impact of many observations
Motivated by the need for practical projections for third-generation instruments and the many-detection era, in this appendix we provide a simple ansatz allowing simple estimates showing how statistical measurement errors on the nuclear equation of state will improve as more observations accumulate. Our argument has five key ingredients. First, we will assume all binaries are equal mass. Second, following Ref. [136] we will assume $\Lambda(m)=\Lambda_{*}(m/m_{*})^{-p}$ with $p=6$, where $\Lambda_{*}$ (the equation of state parameter to be measured with many observations) is the value of the tidal deformability at a reference mass $m_{*}$ (assumed to be at or close to $m_{*}=1.4~{}{\rm M_{\odot}}$). Third, following the discussion in Section III.A of Ref. [93], we will argue the marginal one-dimensional measurement accuracy for the tidal deformability scales as $\sigma_{\Lambda}=\sigma_{*}(m/m_{*})^{-1}$ versus binary mass, adopting a fiducial uncertainty for the faintest source used in our analysis of $\sigma_{*}\simeq 200$; see also, e.g., Fig. 1 in [172]. Fourth, we will assume that each independent observation’s marginal tidal deformability has a log-likelihood of the form
$\Gamma_{**,1}=\Big{(}\frac{\rho}{8}\Big{)}^{2}\sigma_{\lambda}^{-2}\Big{(}\frac{d\Lambda}{d\Lambda_{*}}\Big{)}^{2}=\Big{(}\frac{\rho}{8}\Big{)}^{2}\sigma_{*}^{-2}\Big{(}\frac{m}{m_{*}}\Big{)}^{2-2p}~{}~{}~{}.$ (14)
Fifth, we will assume the signal amplitude $\rho$ scales with mass and luminosity distance as $(m/m_{*})^{5/6}d_{L,*}/d_{L}$ times geometrical angular factors of order unity, i.e.
@@
where in the second line we have replaced the discrete sum by an ensemble average ($\sum_{k}\cdot=N{\left\langle\cdot\right\rangle}$) and for convenience in our ansatz we assume the angular, distance, and mass distributions of the $N$ detected sources are independent. The average ${\left\langle w^{2}\right\rangle}$ of the geometrical factors depends on the assumed network beampattern (and, for realistic heterogeneous networks, their duty cycle), but is well known to be ${\left\langle w^{2}\right\rangle}=2/5$ for the single-interferometer scenario; see Refs. [175, 176, 177] for discussion. Unless otherwise noted, for the purposes of discussion we will assume the single-interferometer beampattern factor, a uniform NS mass distribution between $1~{}{\rm M_{\odot}}$ and $2~{}{\rm M_{\odot}}$, and use ${\left\langle\Big{(}\frac{d_{L,*}}{d_{L}}\Big{)}^{2}\right\rangle}\simeq 1$ as appropriate to a third-generation network. A larger value of this expression corresponds to smaller marginal measurement accuracy $\sigma_{\Lambda_{*,N}}\equiv 1/\sqrt{\Gamma_{**,N}}$ on the fiducial tidal deformability $\Lambda_{*}$.
@@
The analytic estimate provided in Eq. (16) lets us quickly assess the relative importance of different contributions to the integral. As a concrete example, the relative contribution of sources at different distances scales as the relative contributions to ${\left\langle 1/d_{L}^{2}\right\rangle}$, which under the conventional assumption of an event rate comparable to comoving volume has a critical contribution from the weakest sources available to the population average. For a Euclidean distribution of sources out to the maximum detectable distance $d_{L,*}$, within our ansatz we find ${\left\langle(d_{L,*}/d_{L})^{2}\right\rangle}=3$. Weak sources play an even more important role if sources occur more frequently at higher redshift, as expected given the increasing star formation rate with redshift. In the extreme limit of strong redshift dependence and cosmological factors, we expect ${\left\langle(d_{L,*}/d_{L})^{2}\right\rangle}\simeq 1$. The primary limitation of this interpretation is the implicit assumption of a Gaussian joint likelihood, which only applies for loud signals (e.g., comparable to GW170817). Even after restricting our calculation to sources louder than $\rho_{c}$, the detection rate of pertinent sources is only reduced by of order $(8/\rho_{c})^{3}$. For third-generation instruments with $O(10^{5})$ sources, many thousands must still make important contributions, even for the most conservative choices for $\rho_{c}$. Moreover, once in the Gaussian-likelihood regime, our calculation scales much more slowly than the naive Euclidean factor, only as $N{\left\langle(d_{L,*}/d_{L})^{2}\right\rangle}$.
@@
Our expression provides an even stronger conclusion about the relative importance of different parts of the mass distribution: the lowest mass binaries overwhelmingly dominate the mass average ${\left\langle m^{11/3-2p}\right\rangle}\simeq{\left\langle m^{-8}\right\rangle}$ (more precisely = $\frac{1-2^{-(8-2/3)}}{9-2/3}$ for the average for a uniform mass distribution from $1~{}{\rm M_{\odot}}$ to $2~{}{\rm M_{\odot}}$). In retrospect, the lowest mass binaries must dominate: $\Lambda(m)$ is vastly larger at the lowest allowed masses. If any binaries exist at these masses their observations will dominate our constraints on the equation of state because their large tidal effects will be easily differentiated from the noise.
@@
Our simple estimate is comparable to (but calibrated above to be slightly more optimistic than) detailed calculations produced in other studies [178, 104, 172]; for example, ref. [104] argues the 90% credible interval is $\Delta\Lambda_{*}\simeq 910/\sqrt{N}$ and thus that the comparable $68\%$ credible interval would be $520/\sqrt{N}$. The aforementioned measurement accuracy on $\Lambda_{*}$ implies a relative measurement accuracy on $\ln R_{*}$ of $\Lambda_{*}^{-1}|d\ln R_{*}/d\ln\Lambda_{*}|/\sqrt{\Gamma_{**,N}}$. To evaluate this expression and put the result in context, we’ll assume for the purposes of illustration the MPA1 equation of state, which produces for $m_{*}=1.4~{}{\rm M_{\odot}}$ the fiducial pair $(\Lambda_{*},R_{*}/{\rm km})=(513,12.5)$. For simplicity again following Zhao and Lattimer [136], we can evaluate this change-of-variable expression by adopting the same scaling $\Lambda=a(m/R)^{-p}$ assumed above for some constant $a\simeq 0.0085$, so $d\ln R_{*}/d\ln\Lambda_{*}=p$. We therefore estimate the measurement accuracy $\sigma_{R_{*},N}$ in $R_{*}$ produced by stacking $N$ measurements to be
@@
Our analysis in this appendix suggests important trends in tension with previous preliminary investigations. Many groups over several years have argued that only the loudest handful of observations will dominate overall EoS inference [171, 179]. Our analytic argument suggests a comparable and potentially even dominant contribution from more numerous fainter events.
In the discussion in this appendix, we have only provided estimates for the statistical accuracy $\sigma_{\Lambda_{*},N}$ and $\sigma_{R_{*},N}$ for the tidal deformability and radius of a fiducial neutron star, respectively, after combining $N$ events. Given proposed network sensitivity models and reference, these expressions can be immediately translated to the operating time needed to constrain these quantities to any target precision. For example, in terms of the proposed detection rate of useful events (scaled to $10^{5}/{\rm yr}$, assuming all events contribute) we find for the simple estimates discussed here that a target accuracy of $\Delta R_{*}=10~{}{\rm m}$ requires an observing time scaling with event rate and fiducial neutron star radius as
@@ Appendix C More alternate perspectives on GW170817
Figure 17 shows the results of an analysis that incorporates an alternative interpretation of GW170817 performed by Lange et al. (in prep), obtained by comparing the TEOBResumS waveform model to the data. The differences in interpretation reported there propagate to modest differences in the mass-radius posterior distribution using the baseline, broad NS population prior.
@@ Appendix D Hyperpipeline examples
In Section II.7 we presented our iterative strategy for hyperparameter inference. This approach has been used for GW source parameter inference with RIFT as well as the equation of state which are presented in the Section III. Here, we provide illustrations showing how our iterative strategy works within the context of EoS hyperparameter inference.
As an example, Figure 18 shows 90% credible intervals estimated at each stage of one of our analyses, along with (in color) the marginal likelihoods accumulated during the analysis. Initial grid is an initial guess made during this work, however the pipeline starts exploring regions to obtain a more complete picture of the posterior. By the 5th, 6th, and 7th iteration the posterior estimation does not change significantly, leading us to a converged posterior distribution.
@@ Appendix E NICER PSR-J0437
One of the NICER collaboration groups announced the observational constraints for PSR-J0437 [142]. The pulsar has a strong a-priori mass constraint from radio observations [180]. The full posterior data release has not occurred, however, here we apply this pulsar’s separate 90% credible intervals for its mass and radius and can construct an approximate Gaussian posterior to explore its effect on the EoS. The reported Radius is $R=11.36^{+0.95}_{-0.63}~{}{\rm km}$ and $M=1.418\pm 0.037~{}{\rm M_{\odot}}$. For this preliminary application, we employ the constraints as a Gaussian with $R=11.36\pm 0.95~{}{\rm km}$ and $M=1.418\pm 0.037~{}{\rm M_{\odot}}$.
Figure 19 displays the posterior Mass-Radius 90% credible interval shown earlier in Fig. 10, and a new posterior obtained upon applying the mass-radius constraint for PSR-J0437. Again, we notice on this preliminary study, the impact of this additional source on the M-R posterior is very minimal.